# An Efficient Numerical Function Optimization Framework for Constrained Nonlinear Robotic Problems

### v1

@@ Abstract
This paper presents a numerical function optimization framework designed for constrained optimization problems in robotics. The tool is designed with real-time considerations and is suitable for online trajectory and control input optimization problems. The proposed framework does not require any analytical representation of the problem and works with constrained block-box optimization functions. The method combines first-order gradient-based line search algorithms with constraint prioritization through nullspace projections onto constraint Jacobian space. The tool is implemented in C++ and provided online for community use, along with some numerical and robotic example implementations presented in this paper.
@@ 1 Introduction
Numerical optimization (Nocedal and Wright (1999)) is one of the most fundamental backbones of robotics. Numerical optimization methods emerge to address problems that have no analytical solutions and are subject to constraints. Such problems constitute a vast variety of applications, some examples can be listed as mechanical design topology optimization for the best jumping performance (Haldane et al. (2017)), behavioral optimizations to perform certain tasks with the least energy consumption (Mei et al. (2004); Paes et al. (2014)), a control input optimization that also accounts for multiple tasks, limits, contact constraints, and closed kinematic chains (Sovukluk et al. (2023a); Khatib et al. (2022)), motion planning for complex robotic systems (Dai et al. (2014); Sovukluk et al. (2023b); Westervelt et al. (2018); Sovukluk et al. (2025)), apex-to-apex periodic behavior optimization for nonlinear and underactuated systems (Sovukluk et al. (2024)), path planning optimizations for autonomous devices (Gasparetto et al. (2015)), and so on.
@@ 2.5 Inequality Constraint Optimization Subroutine
The inequality constraints are less straightforward to handle than the equality constraints due to two reasons. First, they may not always be active. Second, they are directional. As a result, these constraints should be checked continuously and activated when necessary. Furthermore, if the iteration direction aligns with the inequality constraint’s gradient direction, the constraint should not be activated as any iteration in the given optimal direction would not violate the constraint. The activation conditions are summarized through an example in Fig. 1. A direction check between two vectors can be easily done through the dot operator. Let $\alpha$ be the angle between two arbitrary vectors $\bm{v}_{0}$ and $\bm{v}_{1}$, then
@@
Figure 1: Inequality constraint activation conditions. The left problem never reaches to the activation point. The middle problem starts from the zero crossing point, but the cost function gradient is already in the same direction as the constraint gradient. Hence, the cost function iteration will not exceed the constraint. The right problem also starts from the zero crossing point, and the cost function and constraint gradients are in reverse directions. The constraint is activated. The cost function gradient should be projected onto the nullspace of the constraint gradient, which results in a zero vector in this case. As a result $x$ remains at $x_{0}$. If the constraint was activated in the same way in the middle problem, $x$ again would remain at $x_{0}$, even though the optimal solution does not violate the constraint.
@@ 5 Conclusion
This paper proposes a numerical function optimization framework designed to solve constrained nonlinear robotic problems in real time. The proposed framework does not require any analytical representation of the problem. Hence, it can also be referred as a constrained black-box optimization tool. The method combines first-order gradient-based line search algorithms with constraint prioritization through nullspace projections of constraint Jacobian space. As a result, it has a slower convergence rate per iteration as the Hessian is ignored. On the other hand, the computational cost per iteration is much less as it only requires basic numerical gradient calculations. The capability of the framework is proven through numerous complex numerical and robotic problems ranging from a simple robotic arm to a complex humanoid robot.

### v3

@@ Abstract
This paper presents a numerical function optimization framework designed for constrained optimization problems in robotics. The tool is designed with real-time considerations and is suitable for online trajectory and control input optimization problems. The proposed framework does not require any analytical representation of the problem and works with constrained block-box optimization functions. The method combines first-order gradient-based line search algorithms with constraint prioritization through nullspace projections onto constraint Jacobian space. The tool is implemented in C++ and provided online for community use, along with some numerical and robotic example implementations presented in the end.
@@ 1 Introduction
Numerical optimization (Nocedal and Wright (1999)) is one of the most fundamental backbones of robotics. Numerical optimization methods emerge to address problems that have no analytical solutions and are subject to constraints. Such problems constitute a vast variety of applications, some examples can be listed as mechanical design topology optimization for the best jumping performance (Haldane et al. (2017)), behavioral optimizations to perform certain tasks with the least energy consumption (Mei et al. (2004); Paes et al. (2014)), a control input optimization that also accounts for multiple tasks, limits, contact constraints, and closed kinematic chains (Sovukluk et al. (2023a); Khatib et al. (2022)), motion planning for complex robotic systems (Dai et al. (2014); Sovukluk et al. (2023b); Westervelt et al. (2018)), apex-to-apex periodic behavior optimization for nonlinear and underactuated systems (Sovukluk et al. (2024)), path planning optimizations for autonomous devices (Gasparetto et al. (2015)), and so on.
@@ 2.5 Inequality Constraint Optimization Subroutine
The inequality constraints are less straightforward to handle than the equality constraints due to two reasons. First, they may not always be active. Second, they are directional. As a result, these constraints should be checked continuously and activated when necessary. Furthermore, if the iteration direction aligns with the inequality constraint’s gradient direction, the constraint should not be activated as any iteration in the given optimal direction would not violate the constraint. The activation conditions are summarized through an example in Fig. 1. A direction check between two vectors can be easily done through the dot operator. Let $\alpha\in\mathbb{R}$ be an angle between two arbitrary vectors $\bm{v}_{0}$ and $\bm{v}_{1}$, then
@@
Figure 1: Inequality constraint activation conditions where the red dot, yellow star, and red region represent the initial condition, desired position, and constraint, respectively. The left problem never reaches the activation point. The middle problem starts from the zero crossing point, but the cost function gradient is already in the same direction as the constraint gradient. Hence, the cost function iteration will not exceed the constraint. The right problem also starts from the zero crossing point, and the cost function and constraint gradients are in reverse directions. The constraint is activated. The cost function gradient should be projected onto the nullspace of the constraint gradient, which results in a zero vector in this case. As a result $x$ remains at $x_{0}$.
@@ 5 Conclusion
This paper proposes a numerical function optimization framework designed to solve constrained nonlinear robotic problems in real-time. The proposed framework does not require any analytical representation of the problem and can also be referred as a constrained black-box optimization tool. The method combines first-order gradient-based line search algorithms with constraint prioritization through nullspace projections of constraint Jacobian space. Consequently, it usually has a slower convergence rate per iteration as the Hessian is ignored. On the other hand, the computational cost per iteration is much less as it only requires basic numerical gradient calculations. The capability of the framework is proven through numerous complex numerical and robotic problems ranging from a simple robotic arm to a complex humanoid robot.